3.
1.77e-7 ev Need Help? t Mastar Master H o 0 rt points Previous Answers SercP11 27 P008 The threshold of dark-adapted (scotopic) vision is 3.9 x 10-11 W/m2 at a central wavelength to its maximum diameter of 8.4 mm, how many photons per second enter the eye? 2.59e13X Your response differs significantly from the correct answer. Rework your solution from the bei Need Help? ReadiWa Read It Watch It 0/5 points Previous Answers SerCP1 27 P011 The work function for Platinum is 6.35 ev (a) vert the value of the work function from electron volts to joules
Answer #1

Given data are scotopic vision = 3.9x10-11 W/m2

diameter(d) = 8.4mm => radius =4.2mm=0.0042m

i just taken wavelength as λ-500nm (as WAVELENGTH VALUE SIDE IS CROPPED IN given QUESTION's IMAGE) .so you can substitute any value of LAMBDA in the below equation and get your anser.

Using

E=hf = h(c/\lambda) = (6.63*10^{-34}* 3*10^8)/500*10^{-9} = 3.98x10^{-19} J

for the energy of each photon at this wavelength.

To turn the threshold which is a power/area into energy/sec

we have to multiply by the area of the pupil which is-5 TI. 2 3.14(0.0042)-- 5 53 * 10.

and then the energy/sec = 3.9*10^{-11}*5.53*10^{-5}

2.15 10-15

This then is >2.15 103.98 105,402photons/sec

Therefore the number of photons per second can enter the eye will be 5,402

Note: in your given question "central wavelength" is not given (or cropped) but i just assumpted a value of 500nm,so in place of that you can replace the original value what you have and get your answer.

thank you .. hope will get positive RATING